OpenStudy (anonymous):

H2SO3+Ca(OH)2=???+H2 ((Complete and balance the equation))

OpenStudy (jwheele1):

one sec

OpenStudy (jwheele1):

H2SO3+Ca(OH)2=???+H2 2 H 2 1 S 0 3 O 3 1 Ca 0 2 (OH) 0 2 H 2

OpenStudy (jwheele1):

start with that

OpenStudy (anonymous):

@jwheele1 what? Sorry, I dont understand

OpenStudy (jwheele1):

well all i did was tell you how many of each element or polyatomic ion were on each side. I can help you balance it but i dont know how to fill in the question mark part =(

OpenStudy (anonymous):

oh i see, thanks anyway

OpenStudy (jfraser):

H2SO3 is sulfurous acid, Ca(OH)2 is calium hydroxide, a base. A reaction between an acid and base always produces water and some kind of salt. the H+ from the acid and the OH- from the base form water. The other product will be some combination of calcium (+2) and sulfite (-2). Balancing comes after getting the reactants and products correct first.

OpenStudy (anonymous):

CaSO4.

OpenStudy (jwheele1):

H2SO3+Ca(OH)2=CaSO4+H2 2 H 2 1 S 1 3 O 4 1 Ca 1 2 (OH) 0 2 H 2

OpenStudy (jwheele1):

Now, you can balance

OpenStudy (jwheele1):

That cant be right becaus there is no OH on the right side....or did you break it up somehow?

OpenStudy (jfraser):

Along with the H+ ion, the hydroxide goes into making water. H_2O and H-OH are the same thing

OpenStudy (jfraser):

CaSO3, not CaSO4. calcium sulfITE is the product, not calcium sulfATE

OpenStudy (anonymous):

sorry .

OpenStudy (jwheele1):

okey dokey...like i mentioned earlier...I am NO chemist, lol